OpenStudy (anonymous):

Karen is going to rent a truck for one day. There are two companies she can choose from, and they have the following prices. Company A charges an initial fee of 25 and an additional 60 cents for every mile driven. Company B charges an initial fee of 30 and an additional 50 cents for every mile driven. For what mileages will Company A charge less than Company B? Write your answer as an inequality, using for the number of miles driven.

OpenStudy (pfenn1):

Let x be number of miles driven, and T_A and T_B equal the total price for Companies A and B, respectively.\[T_{A}=25+0.60x\]\[T_B=30+0.50x\]We want to know for what x will the charge for Company A be less than Company B.\[T_A <T_B\]Can you solve this?

OpenStudy (anonymous):

i get it less than

OpenStudy (anonymous):

should i simplify this or leave it as is?

OpenStudy (pfenn1):

You need to simplify and solve for x.

OpenStudy (anonymous):

thanks for not doubting my intelligence i thought thats what i had to do